OpenStudy (anonymous):

What Is The Value Of n So That The Expression X² + 12 x + n Is A Perfect Square Trinomial ? ( 1 Point ) A . 36 B . 36 C . 72 D . 144

Parth (parthkohli):

HINT: \[x^2 + 6x + 6x + n\]

Parth (parthkohli):

\[x(x + 6) + 6(x + \cdots)\]

Parth (parthkohli):

What do you think? @CaritaDeAngel

OpenStudy (anonymous):

I'm Not So Sure Still

Parth (parthkohli):

Do you know that if you have a perfect square polynomial, then you must separate the middle term into the two same numbers.

Parth (parthkohli):

Here, \(12 x = 6x + 6x\).

OpenStudy (skullpatrol):

$$a^2 +2ab+b^2 = (a+b)^2$$ $$x² + 12 x + n$$

Parth (parthkohli):

Now factor the first two terms: \(x^2 + 6x = x(x + 6)\)

OpenStudy (anonymous):

Is It D ?

Parth (parthkohli):

\[(x + a)^2 = x^2 + 2ax + a^2\]\(12x= 2ax\)

OpenStudy (skullpatrol):

Why do both A and B have the same answer @CaritaDeAngel ?

Parth (parthkohli):

How do you think it is D? Can you explain?

OpenStudy (anonymous):

Oops A Is Actually Supposed To Be 6

OpenStudy (anonymous):

Sorry My Mistake I Didn't See That

OpenStudy (skullpatrol):

No problem :)

OpenStudy (anonymous):

I'll Re-Type The Answer Choices Again In Case You Get Confused A . 6 B . 36 C . 72 D . 144

OpenStudy (skullpatrol):